# GL1-02-06-24-2-supplement

### v2

@@ 3 Characterizing Distributional Parallel Trends
The (conditional) distributional parallel trends are said to hold if for every covariate value $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I\subseteq\mathbb{R}$, such that
$0<F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right)<1$ (9)
for all $v\in I$.
We provide two characterizations of the distributional parallel trends within our threshold-crossing framework in Proposition 1 below and Proposition A.1 in Appendix A. The formula in (9) is akin to the identification result of the (unconditional) continuous outcome CiC model of Athey and Imbens (2006), see Theorem 3.1 of their paper. Note that the unconditional distributional parallel trends can be defined similarly if the focus is on the unconditional distribution. Therefore, without considering additional covariates, Definition 1 is graphically illustrated in Figure 1. Here, the counterfactual distribution is constructed by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment). When $I=\mathbb{R}$, as argued in Athey and Imbens (2006), the quantiles of the counterfactual distribution $Y^{\ast}_{11}(0)$ may be constructed by inverting the equality in (9) for each level $q\in[0,1]$:
$F_{Y_{11}^{*}(0)}^{-1}(q)=F_{Y_{01}^{*}(0)}^{-1}\left(F_{Y_{00}^{*}(0)}\left(F_{Y_{10}^{*}(0)}^{-1}(q)\right)\right).$
Note that the class of models implied by the distributional parallel trends above is generally not nested within the parallel trends model set out in (2) of Roth and Sant’Anna (2023). Specifically, denoting $F_{Y_{gt}(0)}$ as the CDF of the potential outcome in the absence of the treatment, the latter relies on the identifying assumption:
$F_{Y_{11}(0)}(y)=F_{Y_{10}(0)}(y)+F_{Y_{01}(0)}(y)-F_{Y_{00}(0)}(y)$ (10)
for all $y\in\{0,1,2\}$. This direct extension of the linear DiD model for the mean to the entire outcome distribution, while intuitive, may give rise to predictions outside the unit interval as the right-hand side (RHS) is not restricted to lie between zero and one. On the other hand, one may attempt to place a similar “parallel trends”-type assumption on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
@@
for all $y\in\mathbb{R}$. However, this may, in turn, incur undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023). [^5] Taking the family of normal distributions as an example, such “parallel trends”-type assumption for the CDFs holds only if $Y_{gt}^{*}(0)$ is either strictly stationary over time within each group or stable across groups in each period (but not a mixture of these cases or others). In our context, this would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time. As can be demonstrated in Proposition A.1 in Appendix A, our distributional parallel trends generally maintain considerable flexibility in how the time trend, the group membership, and the observed heterogeneity affect the counterfactual outcome. In this respect, we argue that this assumption is sensible for outcomes with discrete, ordered responses.
@@
Our first result characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
@@
Suppose Assumption 1 holds. Then, distributional parallel trends defined in Definition 1 hold if and only if
@@
for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
@@
Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that the distributional parallel trends hold if and only if a weighted “parallel trends”-type assumption holds for the functional median (mean) parameters and a proportional “parallel trends”-type holds for the scaling parameters. We note that if the time exclusively affects the median (or mean) parameters, then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD. More specifically, as can be seen in the first part of (11), the change in the median (or average) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors. Here, the weights play the role of correcting the part of the common time trends that cannot be eliminated by the simple mean difference over time within groups. On the other hand, the second part of (11) states that the proportional change in the scaling parameters for the treated group over time, $\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}$, in the absence of treatment, must be identical to the proportional change for the treated group over the same periods, $\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}$. The equality in (11) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Sections 4 and 5.
It is important to recognize that the results of Proposition 1 remain applicable even when strictly monotonic transformations are applied to the latent outcome $Y_{gt}^{\ast}(0)$, a characteristic inherent to the continuous CiC model (see Athey and Imbens, 2006). [^6] This differs from the invariance of parallel trends discussed in Roth and Sant’Anna (2023), which focuses on the average untreated potential outcomes. This aspect is particularly relevant for data with discrete ordered outcomes, which are often re-categorized in different ways to verify the “robustness” of empirical findings. For instance, in a $J+1$-level scenario (ranging from 00 to $J$), it is common to group the first $j_{1}$ (with $0\leq j_{1}<J$) levels into a new level 0, and then levels $j_{1}+1$ to $j_{2}$ (with $j_{1}+1\leq j_{2}<J$) into a new level 1. Such reclassification constitutes, in essence, a strictly monotonic transformation of $Y_{gt}^{\ast}(0)$, thereby preserving the applicability of Proposition 1.777Consider the model with $J+1$ levels: $Y_{gt}(0)=\sum_{j=0}^{J}\mathbb{I}\left[\kappa_{j}<Y^{*}_{gt}(0)\leq\kappa_{j+1}\right]$, where $\kappa_{j}<\kappa_{j+1}$ for all $0\leq j\leq J$, $\kappa_{0}=-\infty$, $\kappa_{J+1}=\infty$, and $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise. Reclassifying the first $j_{1}$ levels as level 0 and then levels $j_{1}+1$ to $j_{2}$ as level 1 implements the monotonic transformation $v\mapsto\frac{v-\kappa_{j_{1}}}{\kappa_{j_{2}}-\kappa_{j_{1}}}$ across all terms in the inequalities, setting $\kappa_{j_{1}}=0$ and $\kappa_{j_{2}}=1$.
Finally, notice that in Appendix A, we establish a second equivalence result to further characterize the class of models that generate the distributional parallel trends under Assumption 1. The result helps to clarify the link between the CiC models considered in Athey and Imbens (2006) and our approach. While being nonnested within the discrete CiC models with covariates in their paper, our model does share the features of an extended class of their continuous CiC model with modified assumptions accommodating additional covariates, see Appendix A for details.
4 Model Without Underreporting
4.1 No Covariates
In this subsection, we establish a point identification result for the counterfactual distribution of interest in the absence of misreporting and other covariates $\mathbf{X}_{gt}$. That is, setting $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ to be constant functions in $\mathbf{X}_{gt}$ (almost surely), the latent consumption $Y_{gt}^{\ast}(0)$ takes the form:
$Y^{*}_{gt}(0)=\eta_{gt}+\lambda_{gt}\varepsilon_{gt},$ (12)
where $\eta_{gt}\in\mathbb{R}$ and $0<\lambda_{gt}<\infty$ by Assumption 1 (see also Yamauchi, 2020). We readily obtain the following impossibility result for the model without covariates and without “false zeros” in the data:

### v3

@@ 3 Parallel Trends for Ordered Outcomes
$Y^{*}_{gt}(0)$ is said to be generated by a CiC model if for every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I_{\mathbf{x}}\subseteq\mathcal{Y}^{*}_{\mathbf{x},11}(0)\cap\mathcal{Y}^{*}_{\mathbf{x},01}(0)$ such that for all $v\in I_{\mathbf{x}}$:
$F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}^{-1}\left(F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})\mid\mathbf{x}\right)=F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right).$ (3)
We provide two characterizations of this CiC model within our threshold-crossing framework in Proposition 1 below and in Proposition LABEL:PROPS2 of the supplementary material, respectively. The formula in (3) closely resembles the unconditional identification result for the CiC model with continuous outcomes by Athey and Imbens (2006), see Theorem 3.1 of their paper. A similar formula can also be defined for the unconditional counterfactual distribution when the focus shifts to an unconditional analysis.
To illustrate Definition 1 without additional covariates, Figure 1 graphically demonstrates the construction of the counterfactual distribution. Specifically, the counterfactual distribution is derived by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment).
Note that (3) has often been compared to other “parallel trends”-type assumptions on the CDFs of the counterfactual distributions (see, e.g., Remark 4 in Roth and Sant’Anna, 2023). One such alternative involves imposing a “parallel trends”-type assumption directly on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
@@
for all $y\in\mathbb{R}$. However, this approach introduces undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023), making it less suitable for modeling discrete outcome. For instance, in the family of normal distributions, such a “parallel trends”-type assumption for the CDFs holds only under restrictive conditions - either $Y_{gt}^{*}(0)$ must exhibit strict stationarity over time within each group, or it must remain stable across groups in each period (but not a mixture of these cases or others). This would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time.
In contrast, our CiC framework for the latent outcome is more flexible and does not impose homogeneity on the model parameters. Moreover, it accommodates standard parametric Probit and Logit DiD models with specified distribution for the error term. [^5] For example, consider a model $Y^{*}(0)=\mathbf{X}^{\prime}\boldsymbol{\eta}+\alpha T+\beta G+\varepsilon$, where $\varepsilon\sim N(0,1)$. It can be readily shown that (3) is satisfied. Given the widespread use of of Probit and Logit DiDs in applied work (e.g. Mayer et al., 2014; Sequeira, 2016; Hoekstra and Sloan, 2022), this flexibility represents a significant advantage of the CiC framework.
@@
Before stating the first identification result, we impose a further (conditional) support condition, which, together with the strict stationarity condition on $\varepsilon_{gt}$, implies $\mathcal{Y}_{\mathbf{x},11}\subseteq\mathcal{Y}_{\mathbf{x},01}$:
Assumption 2.
For every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and and $(g,t)\in\{0,1\}\times\{0,1\}$, $\mathcal{Y}^{*}_{\mathbf{x},11}(0)\subseteq\mathcal{Y}^{*}_{\mathbf{x},10}(0),$ and $\mathrm{int}\left(\mathcal{Y}^{*}_{\mathbf{x},11}(0)\right)$ is nonempty.
The first proposition characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
@@
Suppose Assumptions 1 and 2 hold. Then, $Y^{*}_{gt}(0)$ is generated by a CiC model as defined in Definition 1 if and only if
@@
for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
@@
Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that $Y^{*}_{gt}(0)$ is generated by a CiC model if and only if a weighted “parallel trends”-type assumption holds for the functional mean (median) parameters and a proportional “parallel trends”-type holds for the scaling parameters. We note that if time trends exclusively affect the mean (or median) parameters (so $\lambda_{g0}(\mathbf{x})=\lambda_{g1}(\mathbf{x})$), then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD models. More specifically, as can be seen in the first part of (4), the change in the mean (or median) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors. Here, the weights play the role of correcting the part of the common time trends that cannot be eliminated by the simple mean difference over time within groups. The equality in (4) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Section 4.
Finally, in the supplementary material, we establish another equivalence result to further characterize the class of CiC models in Definition 1. The result helps to clarify the link between the CiC models considered in Athey and Imbens (2006) and our approach.
4 Underreported Outcomes
As emphasized in the introduction, a persistent challenge in analyzing sensitive behaviors, such as the consumption of “social bads”, is the presence of “false zeros” or, more broadly, underreporting. Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019). This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf. Poirier, 1980). To account for this issue, we introduce potential “reported” consumption, denoted by $C(d)$, $d\in\{0,1\}$, which allows for the possibility that $C(d)\neq Y(d)$ with positive probability. The observability rule is then given by $C=D\cdot C(1)+(1-D)\cdot C(0)$.
It is well established that the CiC model with discrete outcomes, even in the absence of misreporting, does not achieve point identification for the counterfactual distribution $F_{Y_{11}(0)}$ (cf. Athey and Imbens, 2006). When underreporting is introduced alongside the discrete nature of the outcome, the identification problem becomes even more complex.
In this section, we address the challenges posed by underreporting in two distinct ways. First, in Subsection 4.1, we derive nonparametric bounds for the counterfactual distribution of the treatment group after treatment by leveraging parallel trends and an exclusion restriction. However, as we will demonstrate, the expressions for the bounds of the counterfactual distribution are only informative when paired with a relatively small upper bound on the cumulative underreporting probability. A different route is taken in Subsection 4.2, where we introduce a semiparametric CiC model for both consumption and reporting decisions, and achieve point identification.
4.1 Partial Identification
In order to derive informative and tight bounds for the potential outcome distribution $F_{Y_{11}(d)}$, $d\in\{0,1\}$, we firstly leverage an instrumental exclusion restriction that does not affect consumption behavior directly.
Assumption 3.
For any $j\in\{0,1\}$ and $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
$\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right).$
An example for an instrumental exclusion $Z_{gt}$ in our context is a “survey cooperation” indicator, which measures respondents’ willingness to cooperate during the survey when consumption behavior is elicited. This indicator is assumed to have no direct effect on consumption $Y_{gt}(0)$, conditional on all other covariates. As is common with instrumental exclusion restrictions, the condition is untestable and the validity of the instrument is subject to the assumptions of the researcher.
Assumption 4.
For any $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
$10<\operatorname{Pr}\left(C_{gt}(d)\leq 0\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<\operatorname{Pr}\left(C_{gt}(d)\leq 1\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<$ 1
Assumption 4 is a non-degeneracy condition on the observed potential outcome $C_{gt}(d)$, $d\in\{0,1\}$, ensuring that all levels $j\in\{0,1,2\}$ lie within the conditional support of reported consumption. Finally, we combine the exclusion restriction from Assumption 3 with a direct bound on the cumulative misreporting probability. The latter becomes necessary because the exclusion restriction alone is insufficient to provide informative bounds for the counterfactual distribution, as demonstrated in Proposition 3 below.
Assumption 5.
For any $j\in\{0,1\}$, $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$:
$\operatorname{Pr}\left(C_{gt}(d)\geq j+1\mid Y_{gt}(d)\leq j,\mathbf{X}_{gt}=\mathbf{x}\right)$ $=0,$
$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x}\right)$ $\leq\alpha,$
where $\alpha\in[0,1]$ is a known constant.
Assumption 5 explicitly characterizes one-sided misreporting (i.e., underreporting). The first part of Assumption 5 states that all kinds of “overreporting probabilities” are zero. Specifically, if the true potential consumption level is less than or equal to $j$, individuals cannot overstate their consumption to any level greater than $j$ in their reports.
The second part of Assumption 5 introduces an explicit upper bound on the “underreporting probabilities”. This bound can reflect the researcher’s institutional knowledge or serve as a sensitivity parameter. For $\alpha=1$, the assumption imposes no restriction on underreporting, whereas $\alpha=0$ prohibits any underreporting. Note that these assumptions generalize the concept of standard misreporting probabilities for binary outcomes (see, e.g., Mondal and Wang, 2024) to cumulative distribution functions in the context of ordered outcomes.
To state the bounds, we introduce an alternative generalized inverse distribution function, denoted by $F^{(-1)}$ (see Athey and Imbens, 2006). Given a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, we define
$F^{(-1)}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty,\infty\},\ q\mapsto F^{(-1)}\left(q\right)$ $:=\sup\left\{y\in\mathcal{Y}\mid q\geq F\left(y\right)\right\},$
where we adopt the convention $\sup\{\emptyset\}=-\infty$ and $\sup\{\mathbb{R}\}=\infty$. We obtain the following result: